OpenStudy (anonymous):

A company makes jewelry. It costs $6,400 to make 10 pieces of jewelry and $11,200 to make 20 pieces of jewelry. Which equation models the cost, C(x), as a linear function of the number of pieces of jewelry, x?

jimthompson5910 (jim_thompson5910):

Slope: m = ( y2 - y1 )/(x2 - x1) m = ( 11200 - 6400 )/(20 - 10) m = ( 4800 )/(10) m = 480 So the slope is m = 480 Now use y = mx+b and then plug in m = 480, x = 10 and y = 6400 y = mx+b y = 480x + b 6400 = 480(10) + b 6400 = 4800 + b 6400 - 4800 = b 1600 = b b = 1600 So the equation is y = 480x+1600 which is also C(x) = 480x+1600

OpenStudy (anonymous):

thanks you jim :D

jimthompson5910 (jim_thompson5910):

You can check the answer If x = 10, then... C(x) = 480x+1600 C(10) = 480(10)+1600 C(10) = 4800+1600 C(10) = 6400 So that works ----------------------- If x = 20, then... C(x) = 480x+1600 C(10) = 480(20)+1600 C(10) = 9600+1600 C(10) = 11200 So that works as well

jimthompson5910 (jim_thompson5910):

yw